OpenStudy (anonymous):

write an equality of the line containing the specified point and perpendicular to the indicated line (0,7), 4x +3=y

OpenStudy (anonymous):

perpendicular means that the slope should be -1/4 (in respect to original slope of 4) so then u have the equation: \[y = -\frac{ 1 }{ 4 }x + c\] then you plug in y for 7 and x for 0 so you get \[7 = -\frac{ 1 }{ 4 }0 + c\] so then c = 7... which makes your function \[y = -\frac{ 1 }{ 4 }x + 7\]